OpenStudy (anonymous):

The function c(x) = 3x − 4 determines how many cupcakes need to be purchased for a birthday party, where x is the number of kids at the party. The party planner uses d(c(x)) to find the amount of money to bring for the cupcake purchase. The function d(x) = 2x + 5. Solve for how much money to bring when there are 7 kids at the party.

OpenStudy (anonymous):

Options : 17 19 25 39

OpenStudy (anonymous):

Basically what I don't understand is the solving, is it like the f(x) g(x) thing?

OpenStudy (anonymous):

@Ab590 @LaceyLeanne @dmezzullo @Country_Boy_Jonathan.

OpenStudy (anonymous):

Ok. If x is the amount of kids and there are 7 kids...Than change x to a 7!

OpenStudy (anonymous):

I'm guessing the c(x) is just a filler?

OpenStudy (anonymous):

I think the answer is either 19 or 25..

OpenStudy (austinl):

We have two individual functions. \(c(x) = 3x − 4\), how many cupcakes to bring \(d(x) = 2x + 5\), how much money for a certain number of cupcakes To determine how much money the party planner needs means that they use the function \(d(c(x))\) Which is, \(d(c(x))=2(3x-4)+5=6x-8+5=6x-3\) If there are seven kids, that means, \(d(c(7))=6(7)-3=\color{blue}{39}\)

OpenStudy (anonymous):

@austinL Oh okay.. thanks so much for explaining it T_T

OpenStudy (austinl):

You are very welcome!